# Non-Invasive Medical Digital Twins using Physics-Informed Self-Supervised Learning

### v1

@@ 3 Digital Twins for Cardiovascular Hemodynamics
In this section, we apply our physics-informed SSL method to perform patient-specific Pressure–Volume (PV) loop analysis in cardiology using echocardiography (cardiac ultrasound). Left ventricular (LV) PV loops illustrate the LV pressure against LV volume at multiple time points during a single cardiac cycle, providing a comprehensive view of a patient’s cardiac function and encoding various hemodynamic parameters like stroke volume, cardiac output, ejection fraction, myocardial contractility, etc., which can be used to diagnose cardiovascular diseases (Burkhoff, 2013). Traditionally, a full characterization of a PV loop requires an invasive PV loop catheterization procedure (Kass et al., 1986). Our objective is to utilize a noninvasive modality (i.e., echocardiography) to identify a digital twin for each patient, allowing us to simulate their individualized PV loops.
Physics-based model. We utilized a lumped-parameter circuit model of cardiac hemodynamics (i.e., blood flow), also known as the Windkessel model (Westerhof et al., 2009), as our underlying physical model $\mathcal{M}(\theta)$. Lumped parameter models are based on the analogy between blood flow in arterial systems and the flow of electric current in a circuit (Figure 1). The model comprises interconnected compartments that are equivalent to elements of an electric circuit. Resistances represent the resistance of blood flow within the blood vessels, accounting for pressure energy losses in the system; capacitors represent the amount of stored stressed blood; diodes describe mitral and aortic valves, and inductances describe the inertial effects due to blood acceleration (Yu, 1998). The blood flow within the system follows the standard Kirchoff’s voltage and current laws. A patient’s digital twin is represented by an electric circuit with a specific parameter instance $\theta$, where $\theta$ corresponds to the components of the electric circuit. The patient’s physiological state $\mathbf{x}$ corresponds to circuit currents and voltages (i.e., pressures and volumes), which enables simulation of the patient’s PV loop in any local heart structure. In this study, we adopt the five-state electric circuit model from (Simaan et al., 2008), illustrated in Figure 2(b).
The Cardiac LV pressure $P_{LV}(t)$ and LV volume $V_{LV}(t)$ are related by the elastance function:
$E(t)=P_{LV}(t)/(V_{LV}(t)-V_{d}),$ (13)
[…]
@@
Noninvasive data. For each patient $i$, we have access to an echocardiogram in the form of ultrasound video data $\mathbf{y}_{i}$ and a measurement of the LV ejection fraction $\bar{\mathbf{x}}_{i}$. Echocardiography is a widely used noninvasive modality for diagnosing cardiovascular diseases; it can directly provide volumes but not pressure measurements. Our goal is to use the echocardiography clip for a patient to predict their entire PV loop through a fully noninvasive process.
@@ 4.1 Echocardiography Data
We test our physics-informed SSL (P-SSL) approach using two echocardiography datasets: EchoNet and CAMUS. The CAMUS dataset (Leclerc et al., 2019) consists of 500 fully annotated cardiac ultrasound videos in 2-chamber view, each with LV volume labels for end systole and diastole ($V_{LV}(t_{ES})$ and $V_{LV}(t_{ED})$). The videos were processed by spatial and temporal padding, with standardized 30-frame videos with a resolution of 256$\times$256 pixels. The EchoNet dataset (Ouyang et al., 2020) comprises 10,030 apical-4-chamber echocardiography videos from routine clinical care at Stanford University Hospital also labeled with $V_{LV}(t_{ES})$ and $V_{LV}(t_{ED})$. These videos were cropped, masked, and down-sampled to a resolution of 112$\times$112 pixels using cubic interpolation. The CAMUS dataset was split into 450 training samples and 50 validation samples. The EchoNet dataset was partitioned into 7,466 training, 1,288 validation, and 1,276 testing samples, respectively.
@@ 4.2 Simulating individualized PV loops via digital twins
We train a 3D-CNN to output a subset of the model parameters $\theta_{i}$. The network consists of four convolutional layers, each followed by max pooling and concluded by a convolutional layer with global average pooling and two fully connected layers. We choose seven of the model parameters listed in Table A.2 to be learned in our model: mitral valve resistance $R_{M}$, aortic valve resistance $R_{A}$, maximum elastance $E_{\max}$, minimum elastance $E_{\min}$, theoretical LV volume at zero pressure $V_{d}$, start LV volume $V_{LV}(0)$, and heart cycle duration $T_{C}$. The remainder of the parameters are fixed to literature values (see Table A.2). We enforce restrictions on the parameters by normalizing values in our activation layer, which scales parameters using a sigmoid function to ensure parameters fall into realistic bounded ranges (as in Table A.2). Parameters $\theta_{i}$ are then passed to a separately trained fixed-weight feed forward neural network $\phi_{\mathcal{M}}$ to output the labeled values of $\bar{\mathbf{x}}_{i}$ from which we construct training loss as the mean square difference between predicted and physician-labeled $V_{LV}(t_{ES})$ and $V_{LV}(t_{ED})$.
@@
The network $\phi_{\mathcal{M}}$ is pretrained by generating 3,840 synthetic data points linearly sampled from realistic parameter ranges mapping parameters $\theta_{i}$ to $V_{LV}(t_{ED})$ and $V_{LV}(t_{ES})$ using a numerical ODE solver (Virtanen et al., 2020). We use two fully connected layers trained to minimize mean square error between predicted and outputted $V_{LV}(t_{ED})$ and $V_{LV}(t_{ES})$.
@@
Once predicting $\theta_{i}$, we can numerically solve for the complete state vector $\mathbf{x}_{i}$ for each patient. The first state in $\mathbf{x}_{i}$ is the patient’s estimated LV volume $V_{LV}$. The elastance function (Eq. (16)) allows us to obtain LV pressure $P_{LV}(t)$ from volume. We plot the two functions $P_{LV}$, $V_{LV}$ against each other over time to obtain the PV loops. This process is deterministic, as the ODEs/elastance function are fixed by patient parameters $\theta_{i}$. We also use $V_{LV}(t_{ED})$ and $V_{LV}(t_{ES})$ to compute ejection fraction (EF) as
$\mbox{EF}=(V_{LV}(t_{ED})-V_{LV}(t_{ES}))/V_{LV}(t_{ED}).$ (14)
We compare parameters and patient PV loops across EF, a routine indicator giving information on cardiac function.
Figure 3: PV loops for patients with normal, high, and low EF.
@@
Validating simulated PV loops using digital twins. Our model has good correlation between labeled and simulated end-systolic and end-diastolic volumes. Our approach (P-SSL) achieve a mean absolute error (MAE) of 7.50% for CAMUS and 5.59% for EchoNet in predicting EF (Table 1), comparable to a baseline that uses a 3D-CNN and fully connected layers to directly approximate $\mathcal{F}^{-1}$ using supervised learning without passing the physics-constrained layers $\phi_{\mathcal{M}}$.
@@
Table 1: Mean Absolute Error (MAE) for true predicted EF by supervised learning and simulated ejection fraction using P-SSL.
Dataset MAE (%)
Supervised 3DCNN P-SSL 3DCNN
CAMUS 7.16 7.50
EchoNet 5.53 5.59
@@
[…]
Our model predicts PV loops that vary by labeled EF demonstrating that the model learns variability in patient physical state from our labels (Figure 3). Predicted parameters also vary across patient EF in our model, and the relative variation in high- and low-EF patient groups were replicated across EchoNet and CAMUS (Figure 4). Notably, the relationship between the parameters and patient groups by EF concur with intuited patterns. Increases in $E_{\max}$ and $E_{\min}$ give an increased differential between pressure and volume, which is more plausible at higher EF. $R_{M}$ and $R_{A}$ are the circuit resistances associated to the mitral valve and aortic valve, respectively. In the mathematical model, increases in these resistances must either increase pressure or decrease blood flow by Ohm’s law. Decreases in either resistance clinical is considered a sign of increased hemodynamic burden, but its diagnostic value for indications including stenosis are controversial and a decrease in resistance is not necessarily predictive of reduce EF (Mascherbauer et al., 2004; Izgi et al., 2007). In mitral regurgitation, valve resistance decreases, but EF is often increased in the beginning of the disease, which could contribute to the inverse pattern for predicted $R_{M}$ (O’Rourke & Crawford, 1984).
4.3 Unsupervised disease diagnosis
The PV loops simulated in patient digital twins can serve as indicators for certain diseases that may not be directly labeled in the dataset. Variations in PV loops clinically enables the diagnosis of diverse cardiac abnormalities, which cannot be predicted using EF and LV volumes alone (Seppelt et al., 2022). We acquired physician labels for Mitral Stenosis (MS) for a subset of the patients in the EchoNet dataset. These labels were not included in the training data, and are not correlated with the EF measurement used to train our model (i.e., AUC of EF in predicting MS was 0.5). A total of 263 of 10,024 patients were identified as having MS. We randomly sampled 100 patients with MS and 100 patients without MS from the available dataset of 10,024 patients. We computed an average PV-loop for both the MS and non-MS groups and compared their respective averages to analyze potential differences in Figure 4(b). The model trend captures a distinctive patterns associated with MS: MS reduces LV pre-load and increases pulmonary venous pressures as illustrated in the right plot of Figure 4(b) (Bastos et al., 2020; Motshabi-Chakane, 2023), which aligns with the expected effect of MS on patients’ PV loop patterns.
4.4 In-silico clinical trials
Using a patient’s digital twin, we can simulate their counterfactual PV loops under a hypothetical intervention in-silico to estimate its effect on the patient or to optimize the treatment parameters (e.g., dosage) (Pappalardo et al., 2019b; Viceconti et al., 2016). Here, we demonstrate the feasibility of in-silico trial simulations under baseline and intervention conditions with the patient-specific digital twin $\mathcal{M}(\theta_{i})$ (Figures 5(a)). In our setting, the lumped-parameter model of (Simaan et al., 2008) is extended by the inclusion of one new state variable and five new parameters to model the addition of an LVAD. Tuning these new parameters and utilizing our model output allows us to simulate the effect of LVAD on an individual patient in a fully non-invasive manner.
The purpose of an LVAD clinically is to improve cardiac output. The intervention has been shown to increase EF in vivo. We demonstrate the same result in-silico using the (Simaan et al., 2008) model (Figure 5). The average EF for the CAMUS and EchoNet populations increase by 17.6% and 18.9%, respectively, with the addition of an LVAD, consistent with reported findings (Dunlay et al., 2013; Drazner, 2022)), with patients having lower pre-LVAD ejection fractions experiencing more significant increases. Figures 5(c) illustrates the distributional change of EF before and after LVAD implantation, indicating a significant right shift in the EF distribution following the procedure.
Furthermore, the influence of rotation speed on the PV loop dynamics is a crucial factor in optimizing the therapeutic effects of LVADs. The rotation speed determines the rate at which blood is drawn from the left ventricle and ejected into the systemic circulation, directly affecting the pressure-volume relationship. Studies, such as those by Simaan et al. (Simaan et al., 2008), have explored the dynamic behavior of LVADs under varying conditions, including different initial rotation speeds. It is essential to note that the choice of rotation speed should be tailored to individual patient needs for optimal therapeutic outcomes. The impact of initial rotation speed on the PV loop, such as changes in end-diastolic and end-systolic volumes, should be carefully considered in the calibration process (Figure 5(b)).
Figure 5: Counterfactual simulations of the LVAD intervention.
5 Related Work
5.1 Inverse problems
Other strategies for inverse problem with partially known forward model include adversarial methods such as CycleGAN for unpaired signals and measurements and AmbientGAN for training from measurements alone with parametric forward model. CycleGAN simultaneously learns the forward and inverse process by minimizes the expected value of the difference between ground truth inputs and outputs and with forward and inverse model composition outputs (Zhu et al., 2017). AmbientGAN learns the inverse model by discriminating between the distribution of the measurements and distribution of the forward model applied to the modeled inputs (Bora et al., 2018). For settings with unlabeled data, (Dubois et al., 2022) pioneers Untrained Physically Informed Neural Networks (UPINNs), training a CNN to reconstruct images by reducing the training error between the original measurements and outputs passed through a known forward model. These set ups differ from our task, which we further compare in Appendix C.
5.2 Physics-Informed Neural Networks (PINNs)
PINN methodologies for inverse problems learn parameters for differential equation models through a training loss defined as the sum of a data loss and a regularization term consisting of the differential equation itself. Several papers have used PINN approaches to study cardiac function, especially blood flow, including papers that use non-invasive medical imaging data. (van Herten et al., 2022) learn parameters and solutions of a model for myocardial perfusion using measurements from MR images. (Kissas et al., 2020) use a PINN approach to calibrate flow simulation model boundary conditions using data extracted from MRIs. (Herrero Martin et al., 2022) study cardiac electrophysiology with PINNs and demonstrate the ability of the method to learn a latent variable. (Sahli Costabal et al., 2020) and (Grandits et al., 2021) use a PINN approach to study cardiac activation. Our set-up differs in that we learn from images rather than through measurements obtained from images, and we predict informative latent parameters directly. Our methodology differs in that the physical model is not integrated into the loss function during training. Instead, we use supervised learning trained on data loss alone to generate a forward model solving the ODE system and transfer this learning (see Appendix C for a further comparison).

### v2

@@ 3 Digital Twins for Cardiovascular Hemodynamics
We apply our physics-informed SSL method to perform patient-specific Pressure–Volume (PV) loop analysis in cardiology using echocardiography (cardiac ultrasound). Left ventricular (LV) PV loops illustrate the LV pressure against LV volume at multiple time points during a single cardiac cycle, providing a comprehensive view of a patient’s cardiac function and encoding various hemodynamic parameters like stroke volume, cardiac output, ejection fraction, myocardial contractility, etc., which can be used to diagnose cardiovascular diseases (7). Traditionally, a full characterization of a PV loop requires an invasive PV loop catheterization procedure (34). Our objective is to utilize a noninvasive modality (i.e., echocardiography) to identify a digital twin for each patient, allowing us to simulate their individualized PV loops.
Physics-based model. We utilized a lumped-parameter circuit model of cardiac hemodynamics (i.e., blood flow), also known as the Windkessel model (78), as our underlying physical model $\mathcal{M}(\theta)$. Lumped parameter models are based on the analogy between blood flow in arterial systems and the flow of electric current in a circuit (Figure 1). The model comprises interconnected compartments that are equivalent to elements of an electric circuit. Resistances represent the resistance of blood flow within the blood vessels, accounting for pressure energy losses in the system; capacitors represent the amount of stored stressed blood; diodes describe mitral and aortic valves, and inductances describe the inertial effects due to blood acceleration (82). The blood flow within the system follows the standard Kirchoff’s voltage and current laws. A patient’s digital twin is represented by an electric circuit with a specific parameter instance $\theta$, where $\theta$ corresponds to the components of the electric circuit. The patient’s physiological state $\mathbf{x}$ corresponds to circuit currents and voltages (i.e., blood flow and pressures), which enables simulation of the patient’s PV loop in any local heart structure. In this study, we adopt the five-state electric circuit model from (65), illustrated in Figure 2(b).
The cardiac LV pressure $P_{LV}(t)$ and LV volume $V_{LV}(t)$ are related by the elastance function, $E(t)=P_{LV}(t)/(V_{LV}(t)-V_{d}),$ which is modeled using a closed form in Eq. (15) in the Appendix. Evolution of $\mathbf{x}$ over time $t$ is governed by a system of ODEs derived from current conservation laws (Appendix Eq. (20)). The model is parameterized by a vector $\theta$ corresponding to bounds on the elastance, circuit resistance, capacitance, inductance and initial states (see Appendix A.2). Given the patient-specific parameters $\theta_{i}$ for the ODE system and elastance function governing cardiac dynamics, we have a unique solution for $\mathbf{x}_{i}$, i.e., $\mathcal{M}$ is injective (Appendix B).
@@
Noninvasive data. For each patient $i$, we have access to an echocardiogram in the form of ultrasound video data $\mathbf{y}_{i}$ and a physician labeled measurement of the LV ejection fraction $\bar{\mathbf{x}}_{i}$. Echocardiography is a widely used noninvasive modality for diagnosing cardiovascular diseases; it can directly provide volumes but not pressure measurements. Our goal is to use the echocardiography clip for a patient to predict their entire PV loop through a fully noninvasive process.
@@ 4.1 Echocardiography Data
We test our physics-informed SSL (Med-Real2Sim) approach using two echocardiography datasets: EchoNet and CAMUS. The CAMUS dataset (43) consists of 500 fully annotated cardiac ultrasound videos in 2-chamber view, each with LV volume labels for end systole and diastole ($V_{LV}(t_{ES})$ and $V_{LV}(t_{ED})$). The videos were processed by spatial and temporal padding, with standardized 30-frame videos with a resolution of 256$\times$256 pixels. The EchoNet dataset (52) comprises 10,030 apical-4-chamber echocardiography videos from routine clinical care at Stanford University Hospital also labeled with $V_{LV}(t_{ES})$ and $V_{LV}(t_{ED})$. These videos were cropped, masked, and down-sampled to a resolution of 112$\times$112 pixels using cubic interpolation. The CAMUS dataset was split into 450 training samples and 50 validation and testing samples. The EchoNet dataset was partitioned into 7,466 training, 1,288 validation, and 1,276 testing samples.
@@ 4.2 Simulating individualized PV loops via digital twins
We train a 3D-CNN implemented using Pytorch to output a subset of the model parameters $\theta_{i}$. The network consists of four convolutional layers, each followed by max pooling and concluded by a convolutional layer with global average pooling and two fully connected layers. We choose seven of the model parameters listed in Table A.2 to be learned in our model: mitral valve resistance $R_{M}$, aortic valve resistance $R_{A}$, maximum elastance $E_{\max}$, minimum elastance $E_{\min}$, theoretical LV volume at zero pressure $V_{d}$, start LV volume $V_{LV}(0)$, and heart cycle duration $T_{C}$. The remainder of the parameters are fixed to literature values (see Table A.2). We enforce restrictions on the parameters by normalizing values in our activation layer, which scales parameters using a sigmoid function to ensure parameters fall into realistic bounded ranges (as in Table A.2). Parameters $\theta_{i}$ are then passed to a separately trained fixed-weight feed forward neural network $\widehat{\phi}_{\mathcal{M}}$ to output the labeled values of $\bar{\mathbf{x}}_{i}$ from which we construct training loss as the mean square difference between predicted and physician-labeled $V_{LV}(t_{ES})$ and $V_{LV}(t_{ED})$.
@@
The network $\widehat{\phi}_{\mathcal{M}}$ is pretrained by generating 3,840 synthetic data points linearly sampled from realistic parameter ranges mapping parameters $\theta_{i}$ to $V_{LV}(t_{ED})$ and $V_{LV}(t_{ES})$ using a numerical ODE solver (72). We use two fully connected layers trained to minimize mean square error between predicted and outputted $V_{LV}(t_{ED})$ and $V_{LV}(t_{ES})$.
@@
Once predicting $\theta_{i}$, we can numerically solve for the complete state vector $\mathbf{x}_{i}$ for each patient. The first state in $\mathbf{x}_{i}$ is the patient’s estimated LV volume $V_{LV}$. The elastance function (Eq. (15)) allows us to obtain LV pressure $P_{LV}(t)$ from volume. We plot the two functions $P_{LV}$, $V_{LV}$ against each other over time to obtain the PV loops. This process is deterministic, as the ODEs/elastance function are fixed by patient parameters $\theta_{i}$. We also use $V_{LV}(t_{ED})$ and $V_{LV}(t_{ES})$ to compute ejection fraction (EF) as $\mbox{EF}=(V_{LV}(t_{ED})-V_{LV}(t_{ES}))/V_{LV}(t_{ED}).$ We compare parameters and patient PV loops across EF, a routine indicator giving information on cardiac function.
@@
Our model has good correlation between labeled and simulated end-systolic and end-diastolic volumes. Our approach, Med-Real2Sim, achieves a mean absolute error (MAE) of 6.81% for CAMUS and 5.40% for EchoNet in predicting EF (Table 1), comparable to a baseline that uses a 3D-CNN and fully connected layers to directly approximate $\mathcal{F}^{-1}$ using supervised learning without passing the physics-constrained layers $\widehat{\phi}_{\mathcal{M}}$.
Dataset MAE (EF)
Supervised 3DCNN Med-Real2Sim
CAMUS 6.58 6.81
EchoNet 5.62 5.40
@@
Table 1: Mean Absolute Error (MAE) for true and predicted EF (in %) by supervised learning and simulated ejection fraction using Med-Real2Sim.
Our model predicts PV loops that vary by labeled EF demonstrating that the model learns variability in patient physical state from our labels (Figure 4(right)). Predicted parameters also vary across patient EF in our model, and the relative variation in high- and low-EF patient groups were replicated across EchoNet and CAMUS (Figure 3). Notably, the relationship between the parameters and patient groups by EF concur with intuited patterns. Increases in $E_{\max}$ and $E_{\min}$ give an increased differential between pressure and volume, which is more plausible at higher EF. $R_{M}$ and $R_{A}$ are the circuit resistances associated to the mitral valve and aortic valve, respectively. In the mathematical model, increases in these resistances must either increase pressure or decrease blood flow by Ohm’s law. Decreases in either resistance clinical is considered a sign of increased hemodynamic burden, but its diagnostic value for indications including stenosis are controversial and a decrease in resistance is not necessarily predictive of reduce EF (45; 31). In mitral regurgitation, valve resistance decreases, but EF is often increased in the beginning of the disease, which could contribute to the inverse pattern for predicted $R_{M}$ (51).
Figure 3: (a) Learned parameters of the digital twin in high- and low-EF patient groups within the EchoNet an CAMUS datasets. (b) Comparison of average (simulated) PV loops in digital twins of non-Mitral Stenosis (MS) patients and MS patients. The plot illustrates differences in simulated hemodynamics in the two groups and agrees with theoretical PV loops for MS patients (47). Depicting of the theoretical PV loop for MS is courtesy of https://cvphysiology.com/heart-disease/hd009a.
4.3 Unsupervised disease diagnosis
The PV loops simulated in patient digital twins can serve as indicators for certain diseases that may not be directly labeled in the dataset. Variations in PV loops clinically enables the diagnosis of diverse cardiac abnormalities, which cannot be predicted using EF and LV volumes alone (64). We acquired physician labels for Mitral Stenosis (MS) for a subset of the patients in the EchoNet dataset. These labels were not included in the training data, and are not correlated with the EF measurement used to train our model (i.e., AUC of EF in predicting MS was 0.5). A total of 263 of 10,024 patients were identified as having MS. We randomly sampled 100 patients with MS and 100 patients without MS from the available dataset of 10,024 patients. We computed an average PV-loop for both the MS and non-MS groups and compared their respective averages to analyze potential differences in Figure 3(b). The model trend captures a distinctive patterns associated with MS: MS reduces LV pre-load and increases pulmonary venous pressures as illustrated in the right plot of Figure 3(b) (3; 47), which aligns with the expected effect of MS on patients’ PV loop patterns.
4.4 In-silico clinical trials
Using a patient’s digital twin, we can simulate their counterfactual PV loops under a hypothetical intervention in-silico to estimate its effect on the patient or to optimize the treatment parameters (e.g., dosage) (55; 71). Here, we demonstrate the feasibility of in-silico trial simulations under baseline and intervention conditions with the patient-specific digital twin $\mathcal{M}(\theta_{i})$ (Figures 4(a)). In our setting, the lumped-parameter model of (65) is extended by the inclusion of one new state variable and five new parameters to model the addition of an LVAD. Tuning these new parameters and utilizing our model output allows us to simulate the effect of LVAD on an individual patient in a fully non-invasive manner.
The LVAD intervention has been shown to increase EF in vivo. We demonstrate the same result in-silico using the (65) model (Figure 4). The average EF for the CAMUS and EchoNet populations increase by 17.6% and 18.9%, respectively, with the addition of an LVAD, consistent with reported findings (19; 17)), with patients having lower pre-LVAD ejection fractions experiencing more significant increases. Figures 4(c) illustrates the distributional change of EF before and after LVAD implantation, indicating a significant right shift in the EF distribution following the procedure.
Furthermore, the influence of rotation speed on the PV loop dynamics is a crucial factor in optimizing the therapeutic effects of LVADs. The rotation speed determines the rate at which blood is drawn from the left ventricle and ejected into the systemic circulation, directly affecting the pressure-volume relationship. Studies, such as those by Simaan et al. (65), have explored the dynamic behavior of LVADs under varying conditions, including different initial rotation speeds. It is essential to note that the choice of rotation speed should be tailored to individual patient needs for optimal therapeutic outcomes. The impact of initial rotation speed on the PV loop, such as changes in end-diastolic and end-systolic volumes, should be carefully considered in the calibration process (Figure 4(b)).
Figure 4: Counterfactual simulations of the LVAD intervention (left): (a),(b),(c). PV loops for patients with normal, high, and low EF (right).
4.5 Comparison of physics-informed approaches
Our physics-informed pretext task is a computationally efficient generative approximation of the dynamics of the map $\mathcal{M}$. We created a second synthetic dataset of 1000 randomly sampled points to test the out of sample prediction of the physics-constrained layers $\widehat{\phi}_{\mathcal{M}}$ and achieve an MAE of 2.30 on EF. The distribution of the loss is uniform across plausible parameter sets suggesting that our network is not subject to poor approximations in extreme cases (Appendix Figure 9). In contrast, other approaches to learn the forward dynamics of differential equation models include PINNs and Neural ODEs. As previously pointed out, these models are trained to learn the dynamics of a single patient, which is incompatible with our goal of learning Eq. (6) with supervised learning.
We compare the ability of our approach to learn the dynamics of the cardiac hemodynamics model for patients to PINN and Neural ODE methods and find it learns just as well on fewer data points (Table 2). Separate PINN and Neural ODE models are trained for each of $N=50$ sets of patient parameters $\theta_{i}$. We generated using numerical methods a synthetic time series of 60 points for each of the five states in the cardiac model: $\mathcal{D}=\{\mathbf{x}_{i}\}_{i=1}^{N}$. From these synthetic data, we trained both a PINN and Neural ODE approach to predict $\mathbf{x}_{i}$ for each of the patients computing average MAE on the 20 best results (Table 2). Recall that our method $\phi_{\mathcal{M}}$ is trained with volume labels (one state) only. We found that only volume labels performed poorly in the Neural ODE settings (Appendix 12). PINNs were implemented in Tensorflow. Neural ODEs were implemented in Pytorch using torchdiffeq (11).
Architecture MAE Training Time (per Epoch) Cost per $\phi_{\mathcal{F}^{-1}}$ Epoch
Model Loss Avg. Pt EF Pt Specific Model (Avg. N=20) Population Model Memory Time
PINN $\mathbf{x}_{i}$ $||f_{\theta_{i}}||$ + $||\mathbf{x}_{i}(t)-\hat{\mathbf{x}_{i}(t)}||$ 7.67 (N=20) 0.16s NA $\mathcal{O}(N\cdot L)$ $\mathcal{O}(N\cdot L)$
Neural ODE $f_{\theta_{i}}$ $||\mathbf{x}_{i}(t)-\hat{\mathbf{x}_{i}(t)}||$ 3.39 (N = 20) 6.12s NA $\mathcal{O}(N)$ $\mathcal{O}(N\cdot L)$
@@ 4.5 Comparison of physics-informed approaches
Table 2: Comparison of physics informed learning. We show the Mean Absolute Error (MAE) on ejection fraction (EF). Here $N$ is the number of patient parameter sets and $L$ is the number of layers in the network ($\widehat{\phi}_{\mathcal{M}}$ or PINN) or the implicit layers i.e. function evaluations of a Neural ODE. Pt = patient.